Suppose the insulating qualities of the wall of a house come mainly from a 4.0-in. layer of brick and an R-19 layer of insulation.What is the total rate of heat loss through such a wall, if its total area is 245 ft2 and the temperature difference across it is 26 F?
Answer #1

\frac{Q}{t}=A\frac{(T_1-T_2)}{R_1+R_2}=\frac{245(26)}{1+19}=318.5